# CAARMA: Class Augmentation with Adversarial Mixup Regularization

### v3

@@ 2 Related-Work
The development of mixup strategies has evolved substantially since Mixup’s original introduction by (Zhang et al., 2018), which generated virtual samples and mixed labels by linearly combining two input samples and their corresponding labels. This pioneering method proved particularly successful in enhancing data diversity and improving generalization in visual classification tasks. Extensions such as ManifoldMix (Verma et al., 2019) applied this concept to hidden layers, while CutMix (Yun et al., 2019) introduced a patch-based approach by blending rectangular sections of images, offering a novel alternative for augmenting training data. Subsequent mixup strategies focused on tailoring data mixing to specific contexts or improving the precision of mixing. Static policies like SmoothMix (Lee et al., 2020), GridMix (Baek et al., 2021), and ResizeMix (Qin et al., 2023) used hand-crafted cutting techniques, while dynamic approaches such as PuzzleMix (Kim et al., 2020) and AlignMix (Venkataramanan et al., 2022) incorporated optimal-transport methods to determine mix regions with greater flexibility. For Vision Transformers, strategies such as TransMix (Chen et al., 2022) and TokenMix (Liu et al., 2022b) focused on leveraging attention mechanisms to refine mixing operations, particularly for transformer architectures (Dosovitskiy et al., 2021). Recent developments have adapted mixup techniques to tasks beyond classification, such as regression. C-Mixup, for instance, applies sample mixing based on label distances, using a symmetric Gaussian kernel to select samples that improve regression performance (Cai et al., 2021). Further enhancing robustness, RC-Mixup integrates C-Mixup with multi-round robust training, creating a feedback loop where C-Mixup helps identify cleaner data, and robust training improves the quality of data for mixing (Liu et al., 2022a). These specialized approaches reveal mixup’s adaptability across various machine learning tasks, enhancing model performance and data resilience. However, it’s important to note that none of these methods involve mixing in the embedding space, which could allow for the creation of entirely new and synthetic class identities,
@@
Recent advancements in synthetic audio generation have emphasized the creation of diverse and high-quality datasets, crucial for training and evaluating audio-based AI models. A notable innovation is ConversaSynth, a framework utilizing large language models (LLMs) to generate synthetic conversational audio across varied persona settings (Gao et al., 2022). This method begins with generating text-based dialogues, which are then rendered into audio using text-to-speech (TTS) systems. The synthetic datasets produced are noted for their realism and topic variety, proving beneficial for tasks such as audio tagging, classification, and multi-speaker speech recognition. These capabilities make ConversaSynth a valuable tool for developing robust, adaptable AI models that can handle diverse audio data and complex conversational contexts. In the domain of speaker verification, SpeechMix introduces a novel method by mixing speech at the waveform level, carefully adjusting ratios to preserve the distinct characteristics of speaker identity (Jindal et al., 2020). However, like many other generative and augmentation techniques, SpeechMix primarily focuses on manipulating known speaker voices rather than generating new identities, thereby limiting its utility for enhancing speaker diversity critical for effective zero-shot learning. In the context of speaker verification, discriminative neural clustering Li et al. (2021) employs clustering techniques to enhance speaker diarization by learning discriminative embeddings, but it does not address class diversity through synthetic class generation as CAARMA does. Synthio employs a unique approach by using text-to-audio (T2A) diffusion models to augment small-scale audio classification datasets (Joassin and Alvarez-Melis, 2022). It enhances compositional diversity and maintains acoustic consistency by aligning T2A-generated synthetic samples with the original dataset using preference optimization. Additionally, exploring style transfer in synthetic audio, recent work by Ueda et al. employs a VITS-based voice conversion model, conditioned on the fundamental frequency (F0), to produce expressive variations from neutral speaker voices (Ueda et al., 2024). This method achieves cross-speaker style transfer in a FastPitch-based TTS system, incorporating a style encoder pre-trained on timbre-perturbed data to prevent speaker leakage. This technique enhances the utility of synthetic data in applications requiring rich stylistic diversity. These developments underline the increasing sophistication of synthetic audio generation techniques, from multi-speaker conversations in ConversaSynth to Synthio’s optimized T2A augmentation for classification, and cross-speaker style transfers with VITS-based models. They collectively demonstrate the power of synthetic data to enrich audio model training by adding diversity and realism. However, these methods still face limitations in generating entirely new speaker identities, which is critical for expanding the range of recognizable voices in speaker verification systems. CAARMA addresses this gap by directly mixing in the embedding space, creating synthetic speakers that enhance zero-shot learning capabilities. CAARMA not only preserves speaker characteristics but also significantly expands the diversity of speaker identities, offering a superior solution for training more robust and adaptable speaker verification systems.

### v4

@@ 2 Related-Work
The development of mixup strategies has evolved substantially since Mixup’s original introduction by (Zhang et al., 2018), which generated virtual samples and mixed labels by linearly combining two input samples and their corresponding labels. This pioneering method proved particularly successful in enhancing data diversity and improving generalization in visual classification tasks. Extensions such as ManifoldMix (Verma et al., 2019) applied this concept to hidden layers, while CutMix (Yun et al., 2019) introduced a patch-based approach by blending rectangular sections of images, offering a novel alternative for augmenting training data. Subsequent mixup strategies focused on tailoring data mixing to specific contexts or improving the precision of mixing. Static policies like SmoothMix (Lee et al., 2020), GridMix (Baek et al., 2021), and ResizeMix (Qin et al., 2020) used hand-crafted cutting techniques, while dynamic approaches such as PuzzleMix (Kim et al., 2020) and AlignMix (Venkataramanan et al., 2022) incorporated optimal-transport methods to determine mix regions with greater flexibility. For Vision Transformers, strategies such as TransMix (Chen et al., 2022) and TokenMix (Liu et al., 2022) focused on leveraging attention mechanisms to refine mixing operations, particularly for transformer architectures (Dosovitskiy et al., 2021). Recent developments have adapted mixup techniques to tasks beyond classification, such as regression. C-Mixup, for instance, applies sample mixing based on label distances, using a symmetric Gaussian kernel to select samples that improve regression performance (Yao et al., 2022). Further enhancing robustness, RC-Mixup integrates C-Mixup with multi-round robust training, creating a feedback loop where C-Mixup helps identify cleaner data, and robust training improves the quality of data for mixing (Hwang et al., 2024). These specialized approaches reveal mixup’s adaptability across various machine learning tasks, enhancing model performance and data resilience. However, it’s important to note that none of these methods involve mixing in the embedding space, which could allow for the creation of entirely new and synthetic class identities,
@@
Recent advancements in synthetic audio generation have emphasized the creation of diverse and high-quality datasets, crucial for training and evaluating audio-based AI models. A notable innovation is ConversaSynth, a framework utilizing large language models (LLMs) to generate synthetic conversational audio across varied persona settings (Kyaw and Chan, 2024). This method begins with generating text-based dialogues, which are then rendered into audio using text-to-speech (TTS) systems. The synthetic datasets produced are noted for their realism and topic variety, proving beneficial for tasks such as audio tagging, classification, and multi-speaker speech recognition. These capabilities make ConversaSynth a valuable tool for developing robust, adaptable AI models that can handle diverse audio data and complex conversational contexts. In the domain of speaker verification, SpeechMix introduces a novel method by mixing speech at the waveform level, carefully adjusting ratios to preserve the distinct characteristics of speaker identity (Jindal et al., 2020). However, like many other generative and augmentation techniques, SpeechMix primarily focuses on manipulating known speaker voices rather than generating new identities, thereby limiting its utility for enhancing speaker diversity critical for effective zero-shot learning. In the context of speaker verification, discriminative neural clustering Li et al. (2021) employs clustering techniques to enhance speaker diarization by learning discriminative embeddings, but it does not address class diversity through synthetic class generation as CAARMA does. Synthio employs a unique approach by using text-to-audio (T2A) diffusion models to augment small-scale audio classification datasets (Ghosh et al., 2026). It enhances compositional diversity and maintains acoustic consistency by aligning T2A-generated synthetic samples with the original dataset using preference optimization. Additionally, exploring style transfer in synthetic audio, recent work by Ueda et al. employs a VITS-based voice conversion model, conditioned on the fundamental frequency (F0), to produce expressive variations from neutral speaker voices (ueda2024expressive). This method achieves cross-speaker style transfer in a FastPitch-based TTS system, incorporating a style encoder pre-trained on timbre-perturbed data to prevent speaker leakage. This technique enhances the utility of synthetic data in applications requiring rich stylistic diversity. These developments underline the increasing sophistication of synthetic audio generation techniques, from multi-speaker conversations in ConversaSynth to Synthio’s optimized T2A augmentation for classification, and cross-speaker style transfers with VITS-based models. They collectively demonstrate the power of synthetic data to enrich audio model training by adding diversity and realism. However, these methods still face limitations in generating entirely new speaker identities, which is critical for expanding the range of recognizable voices in speaker verification systems. CAARMA addresses this gap by directly mixing in the embedding space, creating synthetic speakers that enhance zero-shot learning capabilities. CAARMA not only preserves speaker characteristics but also significantly expands the diversity of speaker identities, offering a superior solution for training more robust and adaptable speaker verification systems.